GlueX TOF Meeting, May 27, 2011
Friday, May 26, 2011
2:00 pm EDT
JLab: CEBAF Center, Room F326
- US and Canada: (866)740-1260 (toll free)
- International: (303)248-0285 (toll call) or look up toll-free number at http://www.readytalk.com/intl
- enter access code followed by the # sign: 3421244#
Talks can be deposited in the directory
/group/halld/www/halldweb/html/talks/2011-2Q on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2011-2Q/ .
- FSU: Sasha Ostrovidov
- JLab: Eugene Chudakov, Mark Ito (chair), Elton Smith, Sascha Somov, Simon Taylor, Ivan Tolstukhin, Beni Zihlmann
Status Report from Paul
Paul Eugenio could not join us, but sent in a written report:
We received the VM2000ma from 3M, and I sent 3 sheets to FIU. We also received the VXS crate, and have swapped it into our test system. We discovered that our test counter PMTs were sagging and separating at the light guide. This was due to using a silicon cookie rather than a glued pmt/light guide joint. We have glued all PMTs now, and Sasha is ready to restart tests.
I have two new students joining our tof effort: Aristedis Tsaris and Hussein Al Goul. They have started getting familiar with our test set up, measuring plateaus, efficiencies, and spectra. They are constructing short paddles and will study single pmt readout and coupling of light from bar to bar, but they will now spend a few weeks at JLab for Hugs.
Status Report from Sasha
Sasha reported on several items.
- The new crate is working well. Now, all modules can be plugged into the single crate with no sag in the power supplies.
- He showed plots on two topics:
- ADC spectrum with Americium source. Paul put together a source/scintillator package that can be used to test PMT's as they arrived. The energy peak is clearly visible.
- Troubleshooting multiple timing peak problem. The troublesome multiple peaks seen previously have been eliminated. They were due to very large pulses coming from discharge of the cathode to wrapping material on the trigger phototubes.
Summary of Goals from Beni
Beni went through the near-term goals based on discussions at previous meetings.
- Repeat resolution measurements with the new components of the set-up, most notably with the absence of spurious peaks in the timing spectra.
- Continue light guide production, and test the various design geometries.
- straight vs. bent
- small vs. large angle of taper
He also talked to Tim and reported that Chuck Hutton will be the designer assigned to the TOF. Beni will talk to Chuck next week and try to get Chuck in touch with the designer at FSU.
Light Guide Material
Mark brought up the plan to buy more light guide material for prototyping. This needs to be picked up again.